OpenStudy (anonymous):

Titus deposited $1,508.49 in a savings account that earns 4.2% simple interest. What will Titus’s account balance be in 8 months?

OpenStudy (anonymous):

In accounting you learn: Interest =Principle x Interest Rate x Time Interest rate, just remember to move it over to the right side of the point Time set dem to total amount like 12 months or 365 days, use num for the time you need 1508.49 (.042) (8/12) = 42.23772 1508.49+42.24=1550.73 In Calculus you learn (1508.49)(1+.042/12)^t=1551.24 Funny how they are about .50 off from each other Most banks actually pay interest with time as 365 days Anyways hope this helps